# Pressure suppresses the density wave order in kagome metal LuNb6Sn6

### v1

@@ Abstract
The density waves that develop in kagome metals ScV6Sn6 and LuNb6Sn6 at low temperature appear to arise from under-filled atomic columns within a V-Sn or Nb-Sn scaffolding. Compressing this network with applied pressure in ScV6Sn6 suppressed the structural transition temperature by constraining atomic rattling and inhibiting the shifts that define the structural modulation. We predicted that the density wave transition in LuNb6Sn6 at 68 K would be suppressed by pressure as well. In this brief study we examine the pressure dependence of the density wave transition by remeasuring resistance vs temperature up to 2.26 GPa. We found the transition temperature is smoothly depressed and disappears around 1.9 GPa. This result not only addresses our prediction, but strengthens the rattling chains origin of structural instabilities in the HfFe6Ge6-type kagome metals.
@@ I Introduction
Researchers have been enamored by the kagome lattice for the last decades due to the exciting behaviors that can arise in materials hosting this structural motif. In metallic systems, the connectivity of atoms on kagome sites generates electronic structures with desirable states like flat bands, van Hove singularities, and Dirac nodes.[1, 2]
Among metallic kagome compounds, structural modulations have emerged as a reoccurring theme. The CsV3Sb5-family,[3, 4] CsCr3Sb5,[5] V3Sb2,[6] La(Ru,Fe)3Si2,[7] Yb0.5Co3Ge3,[8] hexagonal-FeGe,[9, 10, 11] and ScV6Sn6 [12] all develop periodic structural distortions on cooling.
@@
The density wave instability observed below 92 K in ScV6Sn6 has received significant attention due to some of its peculiar features. These include curious transport properties,[13, 14, 15, 16] competing density wave instabilities,[17, 18, 19, 20, 21, 22, 23] and subtle band structure reconstruction.[24, 25, 26, 27, 28, 29]
The structural modulation in ScV6Sn6 has been interpreted to not arise from the Peierls-like Fermi surface nesting common for charge density wave systems.[23, 30, 31, 17, 18, 19, 28] Instead, extra space in the HfFe6Ge6-type structure promotes rattling of Sc-Sn chains[32] and ultimately the development of static order driven by modulating bonding between Sn1 atoms.[33, 19] This explains why the density wave in ScV6Sn6 can be suppressed by pressure[34] and doping by larger rare earths.[32] In short, HfFe6Ge6-type compounds can have rattling instabilities whenever a small atom sits in a sufficiently large scaffolding.
@@
We recently demonstrated this principle through our synthesis and examination of the extensive Nb6Sn6 family of compounds.[35] The rattling of atoms was enhanced with decreasing rare earth size until the a density wave was observed in LuNb6Sn6 at 68 K. This observation reinforced the rattling chain model developed for ScV6Sn6[32] and inspired an important prediction: We hypothesized that the density wave transition temperature of LuNb6Sn6 will be suppressed with pressure. We anticipate that rattling and Sn1-Sn1 bond modulation will be inhibited as the Nb-Sn framework is compressed as seen in ScV6Sn6.
@@
In this paper we present an expedient test of this idea by measuring the temperature-dependent resistance of LuNb6Sn6 under pressure to 2.26 GPa. We clearly observe the transition signature in the resistance data march to lower temperatures as pressure is increased. The density wave in LuNb6Sn6 is suppressed by applied pressure, confirming our prediction. We further compare the pressure evolution of the density wave transition in ScV6Sn6 and Lu-doped ScV6Sn6 reported previously. Our work reaffirms the applicability of the rattling chain model in under-filled network compounds like the HfFe6Ge6-type materials.
@@ II Methods
Single crystal growth of LuNb6Sn6 was performed using the self-flux technique.[35] Elemental reagents were combined in a ratio of 8:2:90 = Lu:Nb:Sn. We utilized Ames Lab Lu metals, Nb powder (Alfa, 99.9%), and Sn shot (Alfa, 99.9%). Reagents were placed into 5 mL Al2O3 (Canfield) crucibles fitted with a catch crucible/porous frit and sealed within fused quartz ampoules with approximately 0.6–0.7 atm of argon cover gas.[36] Samples were heated to 1150 °C and thermalized for 18 h before cooling to 900 °C at a rate of 1 °C/hr. Samples are subsequently centrifuged at 900 °C. Single crystals are small, well-faceted, gray-metallic, hexagonal plates and blocks. The samples are stable in air, water, and common solvents.
For these high-pressure measurements, a 4 mm ID hybrid cell was used with a BeCu outer wall and NiCrAl alloy for an insert. Copper wires and a fiber optic were fixed in place through the sample platform with Stycast 2850 FT epoxy. A hexagonal block-shaped crystal of LuNb6Sn6 was hand polished into a minuscule bar with a 80120 m cross section using sand paper. Contacts to the sample were made with 0.0005 in platinum wire in a conventional 4-contact arrangement held in place with Epotek H20E silver epoxy cured for 30 min at 130 °C. The pressure medium is Daphne 7575 oil which has been shown to remain hydrostatic to pressures above 3 GPa.[37] Transport measurements of the sample were made with a Lake Shore 372 resistance bridge with 3708 pre-amp/scanner using 1 mA and a 5 s averaging time. All data was collected with cooling curves taken at a rate of 0.4 K/min. Accurate temperatures were ensured by fixing a Cernox thermometer to the outside of the pressure cell and allowing for additional exchange gas in the sample space of the PPMS where the experiment took place. Resistance data was averaged over 15 measurements to reduce the noise in the derivative.
@@ III Results
Figure 1a presents the normalized resistance vs temperature data obtained for a series of pressure. At the lowest pressure, resistivity drops sharply on cooling through 64 K. This correponds to the development of the density wave modulation observed at zero pressure at 68 K.[35] In ScV6Sn6 this drop has been tied to an increase in carrier mobility in the density wave phase.[12, 39]
The shape of the resistance curve evolves neatly with increasing pressure (lighter colors). The cliff-like drop at 0.1 GPa evolves into a cusp-shaped peak at 0.54 and 0.78 GPa. This further morphs into a broader step-up on cooling exemplified by the 1.26 GPa curve. The pressure evolution of this resistance feature mimics that observed in ScV6Sn6[34] and (Lu0.944Lu0.054)V6Sn6.[32] Importantly, the high-pressure jump-up feature is most dramatic in LuNb6Sn6.
There is a sharp drop in resistance between 2.8 and 3.7 K. A linear fit to the onset temperature vs pressure yields a slope of -0.410.01 K/GPa with an intercept of 3.740.02 K. This agrees exquisitely with the pressure dependence the superconducting critical temperature of tin metal ().[40] Despite careful polishing, there is a Sn inclusion in our sample.
@@ IV Discussion
We have confirmed our prediction that the density wave phase in LuNb6Sn6 is suppressed by pressure. This supports the rattling chains picture developed for ScV6Sn6[32] we extended to LuNb6Sn6.[35] Compressing the Nb-Sn scaffolding removes the extra space in the structure, inhibiting density wave formation.
@@
Demonstrating broader applicability of the rattling chain model strengthens its value for materials design. Displacement fluctuations can be engineered by under-filling channels in compounds with rigid scaffolding networks. This could to enhance thermoelectric performance and induce density wave transitions. In addition, identifying the critical pressure required to prevent density wave formation will aid computational efforts exploring the criteria for structural instabilities in the HfFe6Ge6-type metals.
@@
The pressure evolution of the transition temperatures in ScV6Sn6 and (Sc0.944Lu0.056)V6Sn6 are plotted fig. 2.[34] Curiously, although the Lu doped material has a significantly reduced transition temperature at zero pressure it shares a comparable critical pressure with the ScV6Sn6. LuNb6Sn6 has a notably lower critical pressure. This difference might be a result of the Nb-Sn framework being more compressible than the V-Sn. This should be addressable with DFT calculations or diffraction experiments under pressure.
@@
Next we turn to a trend common to all three materials, the evolution of the transition signature shape with pressure. In most density wave phases the resistance is higher in the low temperature phase due to a loss of carrier when band reconstruct.[41] In ScV6Sn6, only the Sn1 p-bands show significant reconstruction accounting for a minor reduction in carrier concentration.[23, 24, 19, 33] It was proposed that the reduced resistance in the density wave phase is due to the strong scattering[12, 13, 39] by low energy phonons at temperatures just above the transition.[17, 18, 20] It is likely the same explanation lies behind the step down transition feature observed in LuNb6Sn6 at zero-pressure.
@@
Why then does the transition step difference change sign on increasing pressure in ScV6Sn6 and LuNb6Sn6? First, this might represent a crossover in the relative impact of the change of carrier concentration and mobility as the density wave develops. Alternatively, there may be a change in the nature of the density wave order at higher pressures. For example, applying pressure might favor the competing 1/3 1/3 1/2 density wave[17, 18, 20] over the 1/3 1/3 1/3 observed at zero-pressure in both compounds.[12, 35] These ideas could be tested by low-temperature Raman spectroscopy, optical conductivity, or x-ray scattering under pressure.
@@ V Conclusion
The density wave instabilities of ScV6Sn6 and LuNb6Sn6 are tied extra space provided by undersized atoms in a rigid V-Sn or Nb-Sn network. In the scandium compound, compressing this network with pressure inhibits the displacements that constitute the density wave, reducing the transition temperature. We have demonstrated that density wave order in LuNb6Sn6 is suppressed by pressure as we predicted. This observation emphasizes the value of the rattling chain model for density wave instabilities in filled network compounds.
@@ Acknowledgements.
- Li et al. [2024] Y. Li, Y. Liu, X. Du, S. Wu, W. Zhao, K. Zhai, Y. Hu, S. Zhang, H. Chen, J. Liu, Y. Yang, C. Peng, M. Hashimoto, D. Lu, Z. Liu, Y. Wang, Y. Chen, G. Cao, and L. Yang, Correlated electronic structure and incipient flat bands of the kagome superconductor CsCr3Sb5 10.48550/ARXIV.2406.03740 (2024), arXiv:2406.03740 [cond-mat.str-el] .
@@
- Hu et al. [2023a] H. Hu, Y. Jiang, D. Călugăru, X. Feng, D. Subires, M. G. Vergniory, C. Felser, S. Blanco-Canosa, and B. A. Bernevig, Kagome materials I: SG 191, ScV6Sn6. flat phonon soft modes and unconventional CDW formation: Microscopic and effective theory 10.48550/ARXIV.2305.15469 (2023a), arXiv:2305.15469 [cond-mat.str-el] .
@@
- Ortiz et al. [2024] B. R. Ortiz, W. R. Meier, G. Pokharel, J. Chamorro, F. Yang, S. Mozaffari, A. Thaler, S. J. G. Alvarado, H. Zhang, D. S. Parker, G. D. Samolyuk, J. A. M. Paddison, J. Yan, F. Ye, S. Sarker, S. D. Wilson, H. Miao, D. Mandrus, and M. A. McGuire, Stability frontiers in the kagome metals; the Nb6Sn6 (:Ce-Lu, Y) family and density-wave transition in LuNb6Sn6 10.48550/ARXIV.2411.10635 (2024), arXiv:2411.10635 [cond-mat.str-el] .
@@
- Hu et al. [2023b] T. Hu, H. Pi, S. Xu, L. Yue, Q. Wu, Q. Liu, S. Zhang, R. Li, X. Zhou, J. Yuan, D. Wu, T. Dong, H. Weng, and N. Wang, Optical spectroscopy and band structure calculations of the structural phase transition in the vanadium-based kagome metal ScV6Sn6, Physical Review B 107, 165119 (2023b).

### v2

@@ Abstract
The density waves that develop in kagome metals ScV6Sn6 and LuNb6Sn6 at low temperature appear to arise from under-filled atomic columns within a V-Sn or Nb-Sn scaffolding. Compressing this network with applied pressure in ScV6Sn6 suppressed the structural transition temperature by constraining atomic rattling and inhibiting the shifts that define the structural modulation. We predicted that the density wave transition in LuNb6Sn6 at 68 K would be suppressed by pressure as well. In this brief study, we examine the pressure dependence of the density wave transition by measuring resistance vs temperature up to 2.26 GPa. We found the transition temperature is smoothly depressed and disappears around 1.9 GPa. This result not only addresses our prediction, but strengthens the rattling chains origin of structural instabilities in the HfFe6Ge6-type kagome metals.
@@ I Introduction
Researchers have been enamored by the kagome lattice for the last decades due to the exciting behaviors that can arise in materials hosting this structural motif. In metallic systems, the connectivity of atoms on kagome sites generates electronic structures with desirable states like flat bands, van Hove singularities, and Dirac nodes.[5, 6] Among metallic kagome compounds, structural modulations have emerged as a reoccurring theme. The CsV3Sb5-family,[7, 8] CsCr3Sb5,[9] V3Sb2,[10] La(Ru,Fe)3Si2,[11] Yb0.5Co3Ge3,[12] hexagonal-FeGe,[13, 14, 15] and ScV6Sn6 [16] all develop periodic structural distortions on cooling.
@@
The density wave instability observed below 92 K in ScV6Sn6 has received significant attention due to some of its peculiar features. These include curious transport properties,[17, 18, 19, 20] competing density wave instabilities,[21, 22, 23, 24, 25, 26, 27] and subtle band structure reconstruction.[28, 29, 30, 31, 32, 33] The structural modulation that develops in ScV6Sn6 has characteristics that distinguish it from traditional charge density waves (CDW). In particular, the transition temperature can be suppressed by compressing the lattice with pressure [34] and by filling the lattice by larger rare earth elements.[35]
This is different from usual behavior expected for the lattice volume dependence for CDW systems. Expanding the lattice through isovalent doping and shrinking it with pressure are expected to have opposite impacts on the Fermi-surface details that underlie CDW formation. Therefore, these system often have common trends of transition temperature with lattice volume with both doping and pressure (Fig. 1a).[1, 2, 3, 4]
In fact, there is direct evidence that the modulation in ScV6Sn6 does not arise from the Fermi surface nesting common for CDWs.[27, 36, 37, 21, 22, 23, 32] This motivated the development of a rattling model to understand the structural instability in the material.[35] In short, HfFe6Ge6-type compounds can have rattling instabilities whenever a small atom sits in a sufficiently large scaffolding. In ScV6Sn6 this extra space leads to the formation of Sn-Sn bond modulation that characterize the low-temperature structure.[22, 37]
We recently demonstrated that this rattling instability model extends to the extensive Nb6Sn6 family of compounds.[38] In the compounds with the smallest rare earths, x-ray experiments reveal rattling of rare earth-Sn chains within the Nb-Sn network (Fig. 1b). Crucially, LuNb6Sn6 develops a density wave (DW) order at approximately 68 K with static Sn-Sn bond modulation (Fig. 1d) analogous to that in ScV6Sn6. We demonstrated that replacing lutetium with larger rare earth elements not only inhibits the rattling at high temperatures, but also prevents the development of a DW (Fig. 1e).
@@
This observation reinforced the rattling chain model developed for ScV6Sn6[35] and inspired an important prediction: We hypothesized that the density wave transition temperature of LuNb6Sn6 will be suppressed with pressure (Fig. 1a). We anticipate that rattling and Sn-Sn bond modulation will be inhibited as the Nb-Sn framework is compressed as seen in ScV6Sn6(Fig. 1c).
@@
In this paper we present an expedient test of this idea by measuring the temperature-dependent resistance of LuNb6Sn6 under pressure to 2.26 GPa. We clearly observe that the transition temperature is suppressed as pressure is increased. The density wave in LuNb6Sn6 is suppressed by applied pressure, confirming our prediction. We further compare the pressure evolution of the density wave transition in ScV6Sn6 and Lu-doped ScV6Sn6 reported previously. Our work reaffirms the applicability of the rattling chain model in under-filled network compounds like the HfFe6Ge6-type materials.
@@ II Methods
Single crystal growth of LuNb6Sn6 was performed using the self-flux technique.[38] Elemental reagents were combined in a ratio of 8:2:90 = Lu:Nb:Sn. We utilized Ames Lab Lu metal, Nb powder (Alfa, 99.9%), and Sn shot (Alfa, 99.9%). Reagents were placed into 5 mL Al2O3 (Canfield) crucibles fitted with a catch crucible/porous frit and sealed within fused quartz ampoules with approximately 0.6–0.7 atm of argon cover gas.[39] Samples were heated to 1150 °C and thermalized for 18 h before cooling to 900 °C at a rate of 1 °C/hr. Samples are subsequently centrifuged at 900 °C. Single crystals are small, well-faceted, gray-metallic, hexagonal plates and blocks. The samples are stable in air, water, and common solvents.
For these high-pressure measurements, a 4 mm ID hybrid cell was used with a BeCu outer wall and NiCrAl alloy for an insert. Copper wires and a fiber optic were fixed in place through the sample platform with Stycast 2850 FT epoxy. A hexagonal block-shaped crystal of LuNb6Sn6 was hand polished into a minuscule bar with a 80120 m cross section using sand paper. Contacts to the sample were made with 0.0005 in platinum wire in a conventional 4-contact arrangement held in place with Epotek H20E silver epoxy cured for 30 min at 130 °C. The pressure medium is Daphne 7575 oil which has been shown to remain hydrostatic to pressures above 3 GPa.[40] Transport measurements of the sample were made with a Lake Shore 372 resistance bridge with 3708 pre-amp/scanner using 1 mA and a 5 s averaging time. All data was collected with cooling curves taken at a rate of 0.4 K/min. Accurate temperatures were ensured by fixing a Cernox thermometer to the outside of the pressure cell and allowing for additional exchange gas in the sample space of the PPMS where the experiment took place. Resistance data was averaged over 15 measurements to reduce the noise in the derivative. Raw data is provided in the supplemental materials.[41]
@@ III Results
Figure 2a presents the normalized resistance vs temperature data obtained for a series of pressure. The labeled pressures were estimated at base temperature. At the lowest pressure, resistivity drops sharply on cooling through 64 K. This corresponds to the development of the density wave modulation observed by x-ray diffraction at zero pressure at 68 K.[38] In ScV6Sn6 this drop has been tied to an increase in carrier mobility in the density wave phase.[16, 43]
The shape of the resistance curve evolves neatly with increasing pressure (lighter colors). The cliff-like drop at 0.1 GPa evolves into a cusp-shaped peak at 0.54 and 0.78 GPa. This further morphs into a broader step-up on cooling exemplified by the 1.26 GPa curve. The pressure evolution of this resistance feature mimics that observed in ScV6Sn6[34] and (Sc0.944Lu0.054)V6Sn6.[35] Importantly, the high-pressure jump-up feature is most dramatic in LuNb6Sn6.
In Fig. 2a there is a sharp drop in resistance between 2.8 and 3.7 K. A linear fit to the onset temperature vs pressure yields a slope of -0.410.01 K/GPa with an intercept of 3.740.02 K. This agrees exquisitely with the pressure dependence the superconducting critical temperature of tin metal ().[44] Despite careful polishing, there is a Sn inclusion in our sample.
@@ IV Discussion
We have confirmed our prediction that the density wave phase in LuNb6Sn6 would be suppressed by pressure. This supports the rattling chains picture developed for ScV6Sn6[35] we extended to LuNb6Sn6.[38] Compressing the Nb-Sn scaffolding removes the extra space in the structure, inhibiting DW formation.
It is also possible that compressing the lattice also modifies electronic structure in a way that could disfavor DW development. We suspect these electronic changes are less important based on numerous studies finding that Fermi surface details are not the key driver of DW order in ScV6Sn6.[27, 36, 37, 21, 22, 23, 32] To address this possibility, we propose studying Sc substituted LuNb6Sn6 to further examine the importance of steric effects in DW development. The rattling chains picture would predict that replacing Lu with tiny Sc would exacerbate the under-filling of the scaffolding increasing the DW transitions temperature.
@@
Demonstrating broader applicability of the rattling chain model beyond the V6Sn6 family strengthens its value for materials design. Displacement fluctuations can be engineered by under-filling channels in compounds with rigid scaffolding networks. This could enhance thermoelectric performance and induce density wave transitions. In addition, identifying the critical pressure required to prevent density wave formation will aid computational efforts exploring the criteria for structural instabilities in the HfFe6Ge6-type metals.
@@
The pressure evolution of the transition temperatures in ScV6Sn6 and (Sc0.944Lu0.056)V6Sn6 are plotted fig. 3.[34] Curiously, although the Lu-doped material has a significantly reduced transition temperature at zero pressure, it shares a comparable critical pressure with ScV6Sn6. LuNb6Sn6 has a notably lower critical pressure. This difference might be a result of the Nb-Sn framework being more compressible than the V-Sn. This should be addressable with DFT calculations or diffraction experiments under pressure.
@@
Next we turn to a trend common to all three materials, the evolution of the shape of the resistance transition signature with pressure. In most density wave phases the resistance is higher in the low temperature phase due to a loss of carrier when band reconstruct.[45] In ScV6Sn6, only the Sn1 p-bands show significant reconstruction accounting for a minor reduction in carrier concentration.[27, 28, 23, 46] It was proposed that the reduced resistance in the DW phase is due to sudden loss of strong scattering[16, 17, 43] by low energy phonons associated with the rattling chains.[21, 22, 24] It is likely the same explanation lies behind the step down transition feature observed in LuNb6Sn6 at zero-pressure.
@@
Why then does the transition step difference change sign on increasing pressure in both ScV6Sn6 and LuNb6Sn6? First, this might represent a crossover in the relative impact of the change of carrier concentration and mobility as the density wave develops. Alternatively, there may be a change in the nature of the density wave order at higher pressures. For example, applying pressure might favor the competing 1/3 1/3 1/2 density wave[21, 22, 24] over the 1/3 1/3 1/3 observed at zero-pressure in both compounds.[16, 38] These ideas could be tested by low-temperature Raman spectroscopy, optical conductivity, or x-ray scattering under pressure.
@@ V Conclusion
The density wave instabilities of ScV6Sn6 and LuNb6Sn6 result from the extra space provided by undersized atoms in a rigid V-Sn or Nb-Sn network. In the scandium compound, compressing this network with pressure inhibits the displacements that constitute the density wave, reducing the transition temperature. We predicted that the transition should be suppressed with pressure in LuNb6Sn6 as well. Our temperature dependent resistance measurements reveal confirm this. This observation emphasizes the value of the rattling chain model for understanding the tunability of density wave instabilities in filled network compounds.
@@ Acknowledgements.
- Sacchetti et al. [2007] A. Sacchetti, E. Arcangeletti, A. Perucchi, L. Baldassarre, P. Postorino, S. Lupi, N. Ru, I. R. Fisher, and L. Degiorgi, Pressure dependence of the charge-density-wave gap in rare-earth tritellurides, Physical Review Letters 98, 026401 (2007).
- Luccas et al. [2015] R. F. Luccas, A. Fente, J. Hanko, A. Correa-Orellana, E. Herrera, E. Climent-Pascual, J. Azpeitia, T. Pérez-Castañeda, M. R. Osorio, E. Salas-Colera, N. M. Nemes, F. J. Mompean, M. García-Hernández, J. G. Rodrigo, M. A. Ramos, I. Guillamón, S. Vieira, and H. Suderow, Charge density wave in layered La1-xCexSb2, Physical Review B 92, 235153 (2015).
- Bud’ko et al. [a] S. L. Bud’ko, S. Huyan, P. Herrera-Siklody, and P. C. Canfield, Rapid suppression of charge density wave transition in LaSb2 under pressure, Philosophical Magazine 103, 561 (a).
- Bud’ko et al. [b] S. L. Bud’ko, T. A. Wiener, R. A. Ribeiro, P. C. Canfield, Y. Lee, T. Vogt, and A. H. Lacerda, Effect of pressure and chemical substitutions on the charge-density-wave in LaAgSb2, Physical Review B 73, 184111 (b).
@@
- Li et al. [2025] Y. Li, Y. Liu, X. Du, S. Wu, W. Zhao, K. Zhai, Y. Hu, S. Zhang, H. Chen, J. Liu, Y. Yang, C. Peng, M. Hashimoto, D. Lu, Z. Liu, Y. Wang, Y. Chen, G. Cao, and L. Yang, Electron correlation and incipient flat bands in the Kagome superconductor CsCr3Sb5, Nature Communications 16, 10.1038/s41467-025-58487-x (2025).
@@
- Hu et al. [2025] H. Hu, Y. Jiang, D. Călugăru, X. Feng, D. Subires, M. G. Vergniory, C. Felser, S. Blanco-Canosa, and B. A. Bernevig, Kagome materials I: SG 191, ScV6Sn6. flat phonon soft modes and unconventional charge density wave formation: Microscopic and effective theory, Physical Review B 111, 054113 (2025).
@@
- Ortiz et al. [2025] B. R. Ortiz, W. R. Meier, G. Pokharel, J. Chamorro, F. Yang, S. Mozaffari, A. Thaler, S. J. Gomez Alvarado, H. Zhang, D. S. Parker, G. D. Samolyuk, J. A. M. Paddison, J. Yan, F. Ye, S. Sarker, S. D. Wilson, H. Miao, D. Mandrus, and M. A. McGuire, Stability frontiers in the Kagome metals: The Nb6Sn6 (:Ce-Lu,Y) family and density-wave transition in LuNb6Sn6, Journal of the American Chemical Society 147, 5279 (2025).
@@
- Hu et al. [2023] T. Hu, H. Pi, S. Xu, L. Yue, Q. Wu, Q. Liu, S. Zhang, R. Li, X. Zhou, J. Yuan, D. Wu, T. Dong, H. Weng, and N. Wang, Optical spectroscopy and band structure calculations of the structural phase transition in the vanadium-based kagome metal ScV6Sn6, Physical Review B 107, 165119 (2023).